The amount of energy needed to power a 0.10-kW bulb for one minute would be just sufficient to lift a 1.0-kg object through a vertical distance of
1
Expert's answer
2019-11-29T09:44:17-0500

Let the electrical energy is equal to gravitation potential energy. Then


"Pt=mgh"

So, the corresponding vertical distance


"h=\frac{Pt}{mg}=\frac{100\:\rm W\times 60\: s}{1\:\rm kg\times 9.8\:\rm m/s^2}=612\:\rm m"